OpenStudy (anonymous):

Advanced Algebra w/ Financial Apps help.. I solved the problem but I must be doing something wrong.

OpenStudy (anonymous):

Question: Larry is 32 years old and starting an IRA (individual retirement account). He is going to invest $250 at the beginning of each month. The account is expected to earn 3.5% interest, compounded monthly. How much money, rounded to the nearest dollar, will Larry have in his IRA if he wants to retire at age 58?

OpenStudy (anonymous):

Answers: $177,075 $176,560 $127,316 $126,946

OpenStudy (anonymous):

My work: [FV=PV(1+i)^{nt}\] 250(1+0.035)^12*26 250(1.035)^312 250(45855.25) FV=11,463,814.34 But that's not an answer choice.. what am I doing wrong? :/

OpenStudy (anonymous):

@texaschic101 can you help?

OpenStudy (anonymous):

@$w3G_Godd

OpenStudy (anonymous):

@$w3G_Godd HALPPPP

OpenStudy (anonymous):

ew ok hold on lol

OpenStudy (anonymous):

yeah its hard i know lol. i just dont understand what im doing wrong? i even tried using an online future value calculator.

OpenStudy (anonymous):

PV = 250 FV = future value (unknown) i = annual interest rate = 0.035 i(m) = monthly interest rate = 1.035^(1/12) - 1 v = 1/(1+i(m)) = monthly discount factor n = # of months = 312 FV = 250 * [(1+i(m))^n - 1] / (1 - v) FV = $126,276.64 This is using interest theory / actuarial notation. To use on a BAII or other financial calculator, use these as the inputs for your TVM keys: N = 312 I/Y = i(m) x 100 = 0.287089872 PV = 0 PMT = -250 Then you compute FV.

OpenStudy (anonymous):

does this help :(

OpenStudy (anonymous):

So the future value would be $126,276.64? That isnt one of the answer choices though.. am I just being retarded right now?

OpenStudy (anonymous):

nevermind yes it is. im retarded and i was looking at a different question. @$w3G_Godd THANK YOU

OpenStudy (anonymous):

aha your welcome ;)

OpenStudy (anonymous):

well wait the first part is right the 126 but not the 276? OH GOD WHY IS THIS HAPPENING

OpenStudy (anonymous):

and ur not retarded lol :P

OpenStudy (anonymous):

idk this question is making my brain hurt lol

OpenStudy (anonymous):

Yeah me too lol this whole class makes my brain hurt. :c

OpenStudy (anonymous):

aww that sucks lol i hope u get it hold on let me try one more time

OpenStudy (anonymous):

250 * 1.035 ^ (26 * 12) At age 58, Larry would have approximately $11, 463, 814 (More exact amount: $11, 463, 814.34) dont ask , me lmao

OpenStudy (anonymous):

^ yeah i tried that one from yahoo answers xD

OpenStudy (anonymous):

lmao ya thts where i got it xP